How do I solve 2x-4y=-8 -4x=8y=-16 using the elimination method? I tried to solve it and got the wrong answer. What steps do I take?
Answer by TimothyLamb(4379)

Question:
How do I solve 2x-4y=-8 -4x=8y=-16 using the elimination method? I tried to solve it and got the wrong answer. What steps do I take?
---
this equation contains a typo, please fix it and re-post:
-4x = 8y = -16
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php